OpenStudy (anonymous):

calculates the limit \[\lim_{x \rightarrow 0 } \ln(x+1)/x\]

OpenStudy (zarkon):

you know L'Hospitals rule?

OpenStudy (shubhamsrg):

on the other hand,,you know the shortcut formulla for 1^(infinity) types ?

OpenStudy (anonymous):

yes i know l'hop rule but how can i solve this limit

OpenStudy (zarkon):

find the derivative of the numerator then find the derivative of the denominator

OpenStudy (zarkon):

\[\lim_{x \rightarrow 0 } \frac{\ln(x+1)}{x}=\lim_{x \rightarrow 0 } \frac{\frac{d}{dx}[\ln(x+1)]}{\frac{d}{dx}[x]}\]

OpenStudy (anonymous):

k

OpenStudy (shubhamsrg):

an alternative would have been in cases like (1+f(x))^(g(x)) where f(x) tends to 0 and g(x) tends to infinity, the ans is e^(f(x) * g(x) ) ..this thing can be derived easily.. here, you had log ( 1+x)^(1/x) if you compare,its the same case.. so ans becomes log(e^(x* 1/x) ) = 1